OpenStudy (anonymous):

Classify this system of linear equations. A. inconsistent B. consistent, independent C. consistent, dependent

OpenStudy (anonymous):

u have read matrix?

OpenStudy (anonymous):

1. x+y=2 x-y=2 consistent and independent

OpenStudy (anonymous):

2. x+y=2 2x+2y=4 consistent and dependent

OpenStudy (anonymous):

3. x+y=2 x+y=4 inconsistent

OpenStudy (anonymous):

it's fine?

OpenStudy (anonymous):

10y=5x=35 x=2y-35

OpenStudy (anonymous):

let me explain...then u can find it by own....consistent and independent means unique solution

OpenStudy (anonymous):

consistent and dependent---infinite solution inconsistent---no solution

OpenStudy (anonymous):

10y=5x=35 ?? x=2y-35

OpenStudy (anonymous):

u have wriiten the equation correctly?

OpenStudy (anonymous):

10y=5x+35 x=2y-35

OpenStudy (anonymous):

5x-10y=35 x-2y=-35 now...compare this with above mentioned three cases..and u can answer it by your own...just try it

OpenStudy (anonymous):

simplify it more...x-2y=7 x-2y=-35

OpenStudy (anonymous):

now compare this with above three cases

OpenStudy (anonymous):

does it look like case 2 ?

OpenStudy (anonymous):

consistent dependent?

OpenStudy (anonymous):

no....it's inconsistent

OpenStudy (anonymous):

case 3

OpenStudy (anonymous):

do u know to plot a linear equation?

OpenStudy (anonymous):

on a graph yes

OpenStudy (anonymous):

i think...today only.....i told u to plot a linear equation

OpenStudy (anonymous):

so if u'll plot these two lines then u'll get two parallel lines...which wont intersect i.e. no solution...

OpenStudy (anonymous):

i.e. inconsistent

OpenStudy (anonymous):

oh yeah alright thanks

OpenStudy (anonymous):

anything more?

OpenStudy (anonymous):

no i'll be good